a particle of mass m moving in a circleof radius r under the action of a force which delievers it constant power p and increases its speed. the angular acceleration at time t is proportional to

a  particle of mass m moving in a circleof radius r under the action of a force which delievers it constant power p and increases its speed. the angular acceleration at time t is proportional to
 

Grade:12th pass

1 Answers

Himaja
9 years ago
Let P = power,
     T = torque,
     w = angular velocity,
     I = moment of inertia
P = T×w
P =[ I ×dw/dt]×w
wdw = P/I dt
∫wdw = P/I ∫dt
w2/2= [P/I]t
w = (2Pt/I)1/2
angular velocity, a = dw/dt
                           = (P/2It)1/2
therefore a is directly proportional to t-1/2.